OpenStudy (anonymous):

Find the length of the side of the square 9x^2 + 6x + 1

OpenStudy (lgbasallote):

i assume that is the area?

OpenStudy (anonymous):

Try factoring that.

OpenStudy (lgbasallote):

factor it then get the square root

OpenStudy (anonymous):

Yes the area

OpenStudy (anonymous):

\[\Large s=\sqrt A=\sqrt{9x^2+6x+1}=\sqrt{(\cdots)^2}=\cdots\]

OpenStudy (anonymous):

I'm still not getting it?

OpenStudy (anonymous):

Factor your polynomial like I suggested, and the answer will become obvious.

OpenStudy (anonymous):

(3x+3) (3x-2) ?

OpenStudy (anonymous):

I don't think I am factoring it correctly?

OpenStudy (anonymous):

That's not correct...you can see that by FOILing that expression and seeing that it doesn't match the original. Try again.

OpenStudy (anonymous):

Hint: the two factors will be equal. (This is why it's a square)

OpenStudy (anonymous):

So when you multiple them they equal 6 and when added they equal 1 ?